### v1

@@ Abstract
We consider the parallel-in-time solution of scalar nonlinear conservation laws in one spatial dimension. The equations are discretized in space with a conservative finite-volume method using weighted essentially non-oscillatory (WENO) reconstructions, and in time with high-order explicit Runge-Kutta methods. The solution of the global, discretized space-time problem is sought via a nonlinear iteration that uses a novel linearization strategy in cases of non-differentiable equations. Under certain choices of discretization and algorithmic parameters, the nonlinear iteration coincides with Newton’s method, although, more generally, it is a preconditioned residual correction scheme. At each nonlinear iteration, the linearized problem takes the form of a certain discretization of a linear conservation law over the space-time domain in question. An approximate parallel-in-time solution of the linearized problem is computed with a single multigrid reduction-in-time (MGRIT) iteration. The MGRIT iteration employs a novel coarse-grid operator that is a modified conservative semi-Lagrangian discretization and generalizes those we have developed previously for non-conservative scalar linear hyperbolic problems. Numerical tests are performed for the inviscid Burgers and Buckley–Leverett equations. For many test problems, the solver converges in just a handful of iterations with convergence rate independent of mesh resolution, including problems with (interacting) shocks and rarefactions.
@@ 1 Introduction
Recently in [16, 14, 17] we developed novel MGRIT/Parareal solvers for scalar, linear hyperbolic PDEs that converge substantially more robustly than existing solvers. Most existing approaches use a coarse-grid problem that is based on discretizing directly the underlying PDE, without regard to the fine-grid discretization. In contrast, we use a semi-Lagrangian discretization of the PDE to ensure that coarse-level information propagates correctly along characteristic curves, but, crucially, the discretization is modified with a correction term to approximately account for the truncation error of the fine-grid problem. The end result is a coarse-grid discretization which faithfully matches the fine-grid discretization, leading to fast and robust convergence of the solver. This coarse-level correction technique is based on that devised in [65] to improve performance of classical geometric multigrid on steady state advection-dominated PDEs. In fact, the issue leading to non-robustness in the steady case, an inadequate coarse-grid correction of smooth characteristic components [3, 65], is also responsible, at least in part, for convergence issues of MGRIT/Parareal on advective problems [17].
@@
$\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}=0,\quad(x,t)\in(-1,1)\times(0,T],\quad u(x,0)=u_{0}(x),$ (cons)
with solution $u=u(x,t)$ and flux function $f=f(u)$. For simplicity we consider periodic spatial boundary conditions $u(-1,t)=u(1,t)$, although other boundary conditions could be used. Our numerical tests consider the Burgers and Buckley–Leverett equations with solutions containing interacting shock and rarefaction waves.
@@
In this work we focus on multilevel-in-time methods. We note, however, that some alternative approaches have been developed recently which do not seem to suffer for some hyperbolic problems, at least to the same extent, as MGRIT and Parareal do (when used with naive direct coarse-grid discretizations) [25, 26, 31, 27, 30, 48, 11, 47, 49, 36]. With the exception of [25, 26], these approaches target temporal parallelism by diagonalizing in time, often using discrete-Fourier-transform-like techniques. While some of these methods have been applied to certain quasilinear hyperbolic-like PDEs (wave equations with nonlinear reaction terms [26, 30, 48]; viscous Burgers equation [64]), we are unaware of their application to PDEs of the form (cons). Thus there is a need for effective parallel-in-time methods applied to nonlinear conservation laws, which we address in this paper.
@@
$\frac{\partial e}{\partial t}+\frac{\partial}{\partial x}\big{(}\alpha(x,t)e\big{)}$ $=0,\quad(x,t)\in(-1,1)\times(0,T],\quad e(x,0)=e_{0}(x),$ (cons-lin)
with solution $e(x,t)$, flux function $f(e,x,t)=\alpha(x,t)e$, and periodic spatial boundary conditions $e(-1,t)=e(1,t)$. To solve these linearized problems with MGRIT, we adapt our existing methodology for (non-conservative) linear hyperbolic PDEs described in [16, 14].
@@
The remainder of this paper is organized as follows. Section 2 details the PDE discretizations we use. Section 3 presents the nonlinear iteration scheme to solve the discretized problems, and Section 4 details the linearization procedure. Section 5 develops an MGRIT iteration for approximately solving the linearized problems, and numerical results are then given in Section 6. Conclusions are drawn in Section 7. Supplementary materials are also included which, primarily, describe some additional details of the MGRIT solver. The MATLAB code used to generate the results in this manuscript can be found at https://github.com/okrzysik/pit-nonlinear-hyperbolic-github.
@@ 2 PDE discretization
$\frac{\operatorname{d}\!\bar{u}_{i}}{\operatorname{d}\!t}=-\frac{f\big{(}u(x_{i+1/2})\big{)}-f\big{(}u(x_{i-1/2})\big{)}}{h},\quad\bar{u}_{i}(t):=\frac{1}{h}\int_{{\cal I}_{i}}u(x,t)\operatorname{d}\!x,$ (1)
@@
$\frac{\operatorname{d}\!\bar{u}_{i}}{\operatorname{d}\!t}\approx-\frac{\widehat{f}_{i+1/2}-\widehat{f}_{i-1/2}}{h}=:\big{(}L(\bar{\bm{u}})\big{)}_{i},\quad i=1,\ldots,n_{x}.$ (2)
@@
$\bar{\bm{u}}^{n,1}$ $=F(\bar{\bm{u}}^{n}),$ (4a)
$\bar{\bm{u}}^{n,2}$ $=\frac{3}{4}\bar{\bm{u}}^{n}+\frac{1}{4}F\big{(}\bar{\bm{u}}^{n,1}),$ (4b)
$\bar{\bm{u}}^{n+1}$ $=\frac{1}{3}\bar{\bm{u}}^{n}+\frac{2}{3}F\big{(}\bar{\bm{u}}^{n,2}\big{)}.$ (4c)
@@
$\bar{\bm{u}}^{n+1}=\Phi(\bar{\bm{u}}^{n}),\quad n=0,1,\ldots,n_{t}-2,$ (5)
with $\Phi\colon\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ the time-stepping operator carrying out some ERK scheme. In this work, the time domain $t\in[0,T]$ of (cons) and (cons-lin) is assumed to be discretized with $n_{t}$ points equispaced by a distance $\delta t$.
When 1st- or 3rd-order FV spatial discretizations are used, they are paired with (3) or (2), respectively. We do not present results for discretizations with order of accuracy greater than three; note that our methodology is extensible to higher-order discretizations but we have not rigorously tested it on such problems. In all cases, the constant time-step size $\delta t$ is chosen such that $c_{\max}:=\max_{u\in[\min(u_{0}(x)),\max(u_{0}(x))]}|f^{\prime}(u)|\frac{\delta t}{h}$ is slightly smaller than one. [^1] When sequential time-stepping, for efficiency reasons, one would typically adapt $\delta t$ at each step so as to step close to the CFL limit rather than fix $\delta t$ as a constant for all steps as we do here. We expect that such adaptive time-stepping could be incorporated into our approach by utilising nested iteration methodology, as has been done with MGRIT [22, 23].
@@ 2.1 Numerical flux
$\nu_{i+1/2}=\nu^{(\rm global)}:=\max\limits_{\begin{subarray}{c}w\in[\min(u_{0}(x)),\max(u_{0}(x))]\\ x\in[-1,1]\end{subarray}}|f^{\prime}(w)|.$ (9)
@@
$\nu_{i+1/2}=\nu_{i+1/2}^{\rm(local)}:=\max\limits_{w\in[\min(u_{i+1/2}^{-},u_{i+1/2}^{+}),\max(u_{i+1/2}^{-},u_{i+1/2}^{+})]}|f^{\prime}(w)|.$ (10)
@@
$\widehat{f}\big{(}e_{i+1/2}^{-},e_{i+1/2}^{+})$ $=\frac{1}{2}\left[\big{(}\alpha_{i+1/2}^{-}+\nu_{i+1/2}\big{)}e_{i+1/2}^{-}+\big{(}\alpha_{i+1/2}^{+}-\nu_{i+1/2}\big{)}e_{i+1/2}^{+}\right],$ (11)
$=:\widehat{f}^{\textrm{lin}}\big{(}\alpha_{i+1/2}^{-},e_{i+1/2}^{-},\alpha_{i+1/2}^{+},e_{i+1/2}^{+},\nu_{i+1/2}\big{)}.$ (12)
@@ 2.2 WENO reconstructions
We now outline how the interface reconstructions $u_{i+1/2}^{\pm}$ that enter into the numerical flux (6) are computed. We begin by introducing polynomial reconstruction, for which further details can be found in Section SM1.
Consider the reconstruction of a smooth function $u(x)$ based on its cell averages. At a given $x\in{\cal I}_{i}$, let $q_{i}^{\ell}(x)\approx u(x)$ be a reconstruction polynomial, of degree at most $k-1$, depending on $k\geq 1$ cell averages of $u$ over the cells $S^{\ell}_{i}:=\{{\cal I}_{i-\ell},\ldots,{\cal I}_{i-\ell+(k-1)}\}$, with left-shift $\ell\in\{0,...,k-1\}$. For each cell ${\cal I}_{i}$, there are $k$ such polynomials since there are $k$ possible left-shifted stencils. Polynomial reconstructions at the left and right interfaces of cell ${\cal I}_{i}$ using a reconstruction stencil with left-shift $\ell$ are then denoted by
$u_{i,-1/2}^{\ell}:=q_{i}^{\ell}\big{(}x_{i-1/2}\big{)}\equiv\big{(}\widetilde{R}^{\ell}\bar{\bm{u}}\big{)}_{i},\quad u_{i,+1/2}^{\ell}:=q_{i}^{\ell}\big{(}x_{i+1/2}\big{)}\equiv\big{(}R^{\ell}\bar{\bm{u}}\big{)}_{i}.$ (13)
@@
To obtain high-order reconstructions of order $12k-$ 1, one then takes weighted combinations of the $k$ different $k$th-order reconstructions with different left-shifts $\ell$. Specifically, we write
$u_{i,-1/2}(\bar{\bm{u}}):=\sum\limits_{\ell=0}^{k-1}\widetilde{b}^{\ell}_{i}(\bar{\bm{u}})\big{(}\widetilde{R}^{\ell}\bar{\bm{u}}\big{)}_{i},\quad u_{i,+1/2}(\bar{\bm{u}}):=\sum\limits_{\ell=0}^{k-1}b^{\ell}_{i}(\bar{\bm{u}})\big{(}R^{\ell}\bar{\bm{u}}\big{)}_{i},$ (14)
using weights $\big{\{}\widetilde{b}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$, $\big{\{}{b}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$. Then, the inputs to the flux (6) are $u_{i+1/2}^{-}=u_{i,+1/2}$ and $u_{i+1/2}^{+}=u_{i+1,-1/2}$. There are several choices for $\big{\{}\widetilde{b}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$, $\big{\{}{b}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$ in (14). In the simplest case, they are taken as the so-called optimal linear weights, $\big{\{}\widetilde{d}^{\ell}\big{\}}$, $\big{\{}{d}^{\ell}\big{\}}$, which are independent of $\bar{\bm{u}}$, and of $i$, and are such that the reconstructions (14) are fully $(2k-1)$-order accurate. For $k>1$, however, $\big{\{}\widetilde{d}^{\ell}\big{\}}$, $\big{\{}{d}^{\ell}\big{\}}$ are not suitable when $u(x)$ lacks sufficient regularity over the large stencil $\cup_{\ell=0}^{k-1}S^{\ell}_{i}$ because they lead to spurious oscillations that do not reduce in size as the mesh is refined. Instead, the typical choice for hyperbolic conservation laws is to use so-called weighted essentially non-oscillatory (WENO) weights, $\big{\{}\widetilde{w}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$, $\big{\{}{w}_{i}^{\ell}(\bar{\bm{u}})\big{\}}$. WENO weights are designed such that over stencils where $u$ is smooth they reduce to, as $h\to 0$, the optimal linear weights, and otherwise they adapt to eliminate contributions to the reconstructions from the specific cells where $u(x)$ is non-smooth.

### v2

@@ Abstract
We consider the parallel-in-time solution of scalar nonlinear conservation laws in one spatial dimension. The equations are discretized in space with a conservative finite-volume method using weighted essentially non-oscillatory (WENO) reconstructions, and in time with high-order explicit Runge-Kutta methods. The solution of the global, discretized space-time problem is sought via a nonlinear iteration that uses a novel linearization strategy in cases of non-differentiable equations. Under certain choices of discretization and algorithmic parameters, the nonlinear iteration coincides with Newton’s method, although, more generally, it is a preconditioned residual correction scheme. At each nonlinear iteration, the linearized problem takes the form of a certain discretization of a linear conservation law over the space-time domain in question. An approximate parallel-in-time solution of the linearized problem is computed with a single multigrid reduction-in-time (MGRIT) iteration, however, any other effective parallel-in-time method could be used in its place. The MGRIT iteration employs a novel coarse-grid operator that is a modified conservative semi-Lagrangian discretization and generalizes those we have developed previously for non-conservative scalar linear hyperbolic problems. Numerical tests are performed for the inviscid Burgers and Buckley–Leverett equations. For many test problems, the solver converges in just a handful of iterations with convergence rate independent of mesh resolution, including problems with (interacting) shocks and rarefactions.
@@ 1 Introduction
Recently in [15, 14, 16] we developed novel MGRIT/Parareal solvers for scalar, linear hyperbolic PDEs that converge substantially more robustly than existing solvers. Most existing approaches use a coarse-grid problem that is based on discretizing directly the underlying PDE, without regard to the fine-grid discretization. In contrast, we use a semi-Lagrangian discretization of the PDE to ensure that coarse-level information propagates correctly along characteristic curves, but, crucially, the discretization is modified with a correction term to approximately account for the truncation error of the fine-grid problem. The end result is a coarse-grid discretization which more faithfully matches the fine-grid discretization, leading to fast and robust convergence of the solver. This coarse-level correction technique is based on that devised in [65] to improve performance of classical geometric multigrid on steady state advection-dominated PDEs. In fact, the issue leading to non-robustness in the steady case, an inadequate coarse-grid correction of smooth characteristic components [3, 65], is also responsible, at least in part, for convergence issues of MGRIT/Parareal on advective problems [16].
@@
(cons) $\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}=0,\quad(x,t)\in(-1,1)\times(0,T],\quad u(x,0)=u_{0}(x),$
with solution $u=u(x,t)$ and flux function $f=f(u)$. For simplicity we consider periodic spatial boundary conditions u(1,t)=u(1,t)u(-1,t)=u(1,t), although other boundary conditions could be used. Our numerical tests consider the Burgers and Buckley–Leverett equations with solutions containing interacting shock and rarefaction waves.
@@
In this work we focus on multilevel-in-time methods, namely MGRIT. We note, however, that some alternative approaches have been developed recently which do not seem to suffer for some hyperbolic problems, at least to the same extent, as MGRIT and Parareal do (when used with naive direct coarse-grid discretizations) [24, 25, 30, 26, 29, 48, 11, 47, 49, 35]. With the exception of [24, 25], these approaches target temporal parallelism by diagonalizing in time, often using discrete-Fourier-transform-like techniques. While some of these methods have been applied to certain quasilinear hyperbolic-like PDEs (wave equations with nonlinear reaction terms [25, 29, 48]; viscous Burgers equation [64]), we are unaware of their application to PDEs of the form (cons). Thus there is a need for effective parallel-in-time methods applied to nonlinear conservation laws, which we address in this paper.
@@
(cons-lin) $\frac{\partial e}{\partial t}+\frac{\partial}{\partial x}\big(\alpha(x,t)e\big)$ $=0,\quad(x,t)\in(-1,1)\times(0,T],\quad e(x,0)=e_{0}(x),$
with solution $e(x,t)$, flux function $f(e,x,t)=\alpha(x,t)e$, and periodic spatial boundary conditions e(1,t)=e(1,t)e(-1,t)=e(1,t). To solve these linearized problems with MGRIT, we adapt our existing methodology for (non-conservative) linear hyperbolic PDEs described in [15, 14]. We emphasize, however, that any other parallel-in-time method effective on linearized problems of the form (cons-lin) could be used in place of MGRIT.
@@
The remainder of this paper is organized as follows. Section 2 details the PDE discretizations we use. Section 3 presents the nonlinear iteration scheme to solve the discretized problems, and Section 4 details the linearization procedure. Section 5 develops an MGRIT iteration for approximately solving the linearized problems, and numerical results are then given in Section 6. Conclusions are drawn in Section 7. Supplementary materials are also included which, primarily, describe some additional details of the MGRIT solver. The MATLAB code used to generate the results in this manuscript can be found at https://github.com/okrzysik/pit-nonlinear-hyperbolic (v1.1.0).
@@ 2 PDE discretization
(1) $\frac{\operatorname{d}\!\bar{u}_{i}}{\operatorname{d}\!t}=-\frac{f\big(u(x_{i+1/2})\big)-f\big(u(x_{i-1/2})\big)}{h},\quad\bar{u}_{i}(t):=\frac{1}{h}\int_{{\cal I}_{i}}u(x,t)\operatorname{d}\!x,$
@@
(2) $\frac{\operatorname{d}\!\bar{u}_{i}}{\operatorname{d}\!t}\approx-\frac{\widehat{f}_{i+1/2}-\widehat{f}_{i-1/2}}{h}=:\big(L(\bar{\bm{u}})\big)_{i},\quad i=1,\ldots,n_{x}.$
@@
(4a) $\bar{\bm{u}}^{n,1}$ $=F(\bar{\bm{u}}^{n}),$
(4b) $\bar{\bm{u}}^{n,2}$ $=\frac{3}{4}\bar{\bm{u}}^{n}+\frac{1}{4}F\big(\bar{\bm{u}}^{n,1}),$
(4c) $\bar{\bm{u}}^{n+1}$ $=\frac{1}{3}\bar{\bm{u}}^{n}+\frac{2}{3}F\big(\bar{\bm{u}}^{n,2}\big).$
@@
(5) $\bar{\bm{u}}^{n+1}=\Phi(\bar{\bm{u}}^{n}),\quad n=0,1,\ldots,n_{t}-2,$
with $\Phi\colon\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ the time-stepping operator carrying out some ERK scheme. In either case, the time discretization is “one step,” meaning that the solution at $t_{n+1}$ is computed only from the solution at the previous time step, $t_{n}$. In this work, the time domain $t\in[0,T]$ of (cons) and (cons-lin) is assumed to be discretized with $n_{t}$ points equispaced by a distance $\delta t$.
When 1st- or 3rd-order FV spatial discretizations are used, they are paired with (3) or (4), respectively. We do not present results for discretizations with order of accuracy greater than three; note that our methodology is extensible to higher-order discretizations but we have not rigorously tested it on such problems. In all cases, the constant time-step size $\delta t$ is chosen such that $c_{\max}:=\max_{u\in[\min(u_{0}(x)),\max(u_{0}(x))]}|f^{\prime}(u)|\frac{\delta t}{h}$ is slightly smaller than one. [^1] When sequential time-stepping, for efficiency reasons, one would typically adapt $\delta t$ at each step so as to step close to the CFL limit rather than fix $\delta t$ as a constant for all steps as we do here. We expect that such adaptive time-stepping could be incorporated into our approach by utilising nested iteration methodology, as has been done with MGRIT for diffusion problems [21, 22].
@@ 2.1 Numerical flux
(7) $\nu_{i+1/2}=\nu^{(\rm global)}:=\max\limits_{\begin{subarray}{c}w\in[\min(u_{0}(x)),\max(u_{0}(x))]\\ x\in[-1,1]\end{subarray}}|f^{\prime}(w)|.$
@@
(8) $\nu_{i+1/2}=\nu_{i+1/2}^{\rm(local)}:=\max\limits_{w\in[\min(u_{i+1/2}^{-},u_{i+1/2}^{+}),\max(u_{i+1/2}^{-},u_{i+1/2}^{+})]}|f^{\prime}(w)|.$
@@
(9) $\widehat{f}\big(e_{i+1/2}^{-},e_{i+1/2}^{+})$ $=\frac{1}{2}\left[\big(\alpha_{i+1/2}^{-}+\nu_{i+1/2}\big)e_{i+1/2}^{-}+\big(\alpha_{i+1/2}^{+}-\nu_{i+1/2}\big)e_{i+1/2}^{+}\right],$
(10) $=:\widehat{f}^{\textrm{lin}}\big(\alpha_{i+1/2}^{-},e_{i+1/2}^{-},\alpha_{i+1/2}^{+},e_{i+1/2}^{+},\nu_{i+1/2}\big).$
@@ 2.2 WENO reconstructions
We now outline how the interface reconstructions $u_{i+1/2}^{\pm}$ that enter into the numerical flux (6) are computed. We begin by introducing polynomial reconstruction, for which further details can be found in Supplementary Materials Section SM1.
Consider the reconstruction of a smooth function $u(x)$ based on its cell averages. At a given $x\in{\cal I}_{i}$, let $q_{i}^{\ell}(x)\approx u(x)$ be a reconstruction polynomial, of degree at most k1k-1, depending on $k\geq 1$ cell averages of $u$ over the cells $S^{\ell}_{i}:=\{{\cal I}_{i-\ell},\ldots,{\cal I}_{i-\ell+(k-1)}\}$, with left-shift $\ell\in\{0,...,k-1\}$. For each cell ${\cal I}_{i}$, there are $k$ such polynomials since there are $k$ possible left-shifted stencils. Polynomial reconstructions at the left and right interfaces of cell ${\cal I}_{i}$ using a reconstruction stencil with left-shift $\ell$ are then denoted by
(11) $u_{i,-1/2}^{\ell}:=q_{i}^{\ell}\big(x_{i-1/2}\big)\equiv\big(\widetilde{R}^{\ell}\bar{\bm{u}}\big)_{i},\quad u_{i,+1/2}^{\ell}:=q_{i}^{\ell}\big(x_{i+1/2}\big)\equiv\big(R^{\ell}\bar{\bm{u}}\big)_{i}.$
@@
To obtain high-order reconstructions of order 2k12k-1, one then takes weighted combinations of the $k$ different $k$th-order reconstructions with different left-shifts $\ell$. Specifically, we write
(12) $u_{i,-1/2}(\bar{\bm{u}}):=\sum\limits_{\ell=0}^{k-1}\widetilde{b}^{\ell}_{i}(\bar{\bm{u}})\big(\widetilde{R}^{\ell}\bar{\bm{u}}\big)_{i},\quad u_{i,+1/2}(\bar{\bm{u}}):=\sum\limits_{\ell=0}^{k-1}b^{\ell}_{i}(\bar{\bm{u}})\big(R^{\ell}\bar{\bm{u}}\big)_{i},$
using weights $\big\{\widetilde{b}_{i}^{\ell}(\bar{\bm{u}})\big\}$, $\big\{{b}_{i}^{\ell}(\bar{\bm{u}})\big\}$. Then, the inputs to the flux (6) are $u_{i+1/2}^{-}=u_{i,+1/2}$ and $u_{i+1/2}^{+}=u_{i+1,-1/2}$. There are several choices for $\big\{\widetilde{b}_{i}^{\ell}(\bar{\bm{u}})\big\}$, $\big\{{b}_{i}^{\ell}(\bar{\bm{u}})\big\}$ in (12). In the simplest case, they are taken as the so-called optimal linear weights, $\big\{\widetilde{d}^{\ell}\big\}$, $\big\{{d}^{\ell}\big\}$, which are independent of $\bar{\bm{u}}$, and of $i$, and are such that the reconstructions (12) are fully (2k1)(2k-1)-order accurate. For $k>1$, however, $\big\{\widetilde{d}^{\ell}\big\}$, $\big\{{d}^{\ell}\big\}$ are not suitable when $u(x)$ lacks sufficient regularity over the large stencil $\cup_{\ell=0}^{k-1}S^{\ell}_{i}$ because they lead to spurious oscillations that do not reduce in size as the mesh is refined. Instead, the typical choice for hyperbolic conservation laws is to use so-called weighted essentially non-oscillatory (WENO) weights, $\big\{\widetilde{w}_{i}^{\ell}(\bar{\bm{u}})\big\}$, $\big\{{w}_{i}^{\ell}(\bar{\bm{u}})\big\}$. WENO weights are designed such that over stencils where $u$ is smooth they reduce to, as $h\to 0$, the optimal linear weights, and otherwise they adapt to eliminate contributions to the reconstructions from the specific cells where $u(x)$ is non-smooth.